### v1

# A Graph Neural Network Approach to Dispersed Systems

Abstract
We present a Graph Neural Network (GNN) that accurately simulates a multidisperse suspension of interacting spherical particles. Our machine learning framework is build upon the recent work of Sanchez-Gonzalez et al. ICML, PMLR, 119, 8459–8468 (2020) on graph network simulators, and efficiently learns the intricate dynamics of the interacting particles. Nodes and edges of the GNN correspond to, respectively, the particles with their individual properties/data (e.g., radius, position, velocity) and the pair interactions between the particles (e.g., electrostatics, hydrodynamics). A key contribution of our work is to account for the finite dimensions of the particles and their impact on the system dynamics. We test our GNN against an exemplary case study of a multidisperse mixture of two-dimensional spheres sedimenting under gravity in a liquid and interacting with each other by a Lennard-Jones potential. The present GNN framework offers a fast and accurate method for the theoretical study of complex physical systems such as field-induced behavior of colloidal suspensions and ionic liquids. Our implementation of the GNN is available on GitHub at github.com/rfjd/GNS-DispersedSystems.
I Introduction
Accurate and efficient simulation of suspensions and emulsions (e.g., colloidal dispersions [1]) is a key component in better understanding the underlying physics of such systems, as well as leveraging them for practical applications spanning from engineering to biology such as cargo-loading and targeted drug-delivery by micro/nano-robots [2, 3, 4], field-induced self-assembly [5, 6], and electrophoretic deposition [7, 8, 9]. Such problems are governed by rather complex, and at times poorly understood, physics. As a result, it is often difficult to develop a comprehensive mathematical model without invoking unrealistic assumptions. Furthermore, even the reduced, approximate, models are often impossible to solve accurately. An alternative approach that circumvents these challenges is to use Machine Learning (ML) to develop a data-driven model.
In a recent seminal work, Sanchez-Gonzalez et al. [10] used Graph Neural Networks (GNNs) [11, 12, 13] to simulate the complex physics of fluid dynamics in a wide range of conditions. The corresponding machine learning framework, referred to as a Graph Network Simulator (GNS), treats the fluid medium (along with the immersed solid objects) as an ensemble of point particles interacting with each other. As such, GNS is similar to particle-based methods such as smoothed particle hydrodynamics [14], position based dynamics [15], and material point method [16].
Figure 1: Schematic diagram of the problem: a multidisperse suspension of interacting spherical particles in a rectangular box of liquid. A target particle $i$, with current position $\boldsymbol{x}_{i}$ and velocity $\boldsymbol{v}_{i}$, moves and interacts with its neighbors.
We extend the GNS [10, 17] by accounting for the finite size of the particles in a fluid suspension. Our methodology acts as an accurate, efficient, surrogate to the often time-consuming, conventional, computational methods such as computational fluid dynamics (CFD). We represent the immersed particles as nodes of a graph, along with their important dynamic and intrinsic properties such as position, velocity, and radius. The interaction between the particles is modeled as edges of the graph; i.e, the pair-interaction between the particles is represented as a message-passing process between the nodes of the graph.
We evaluate the accuracy of our GNN in an exemplary case-study of a multidisperse suspension of particles sedimenting under gravity in a box of fluid. A comparison between the ground-truth data and predictions of the trained GNN indicates a great agreement. The trained GNN accurately and efficiently captures the most intricate characteristics of the system behavior such as individual particle trajectories, clusters, collisions, and boundaries. Our implementation of the GNN is publicly available on GitHub at github.com/rfjd/GNS-DispersedSystems.
II Problem Statement & GNN
II.1 Multidisperse System of Interacting Spheres
We consider $N$ spherical particles (or droplets) with finite radii $a_{i},\;i=1,2,\cdots,N$, interacting and moving in a box of liquid (see Fig. 1). Depending on the size and charge of the particles, and the type of the liquid, various interactions such as electrostatics, hydrodynamics, steric, and van der Waals forces can be present. Also, the particles experience a drag force from the liquid medium, and may move due to spatially and temporally uniform or nonuniform external forces such as those from a gravitational field or an imposed electric field.
The training data contains trajectories of all particles with a time-step $\Delta t$. Such training data can be obtained from experiments or computer simulations. The objective is to train a GNN that understands the dynamics of such system and can accurately predict the trajectories of all particles. In case of computer simulations, the idea is that the trained GNN (i.e., our data-driven model) is much faster than the physics-based simulator. As shown in Fig. 1, we denote the position and velocity of particle $i$ with $\boldsymbol{x}_{i}$ and $\boldsymbol{v}_{i}$, respectively. Furthermore, $\boldsymbol{F}_{ij}$ is the force on particle $i$ due to the neighboring particle $j$. Of course, we assume that we do not have any information about the nature of this force; instead we see the outcome of the interaction through the trajectories of the particles. Following Sanchez-Gonzalez et al. [10], we use a first-order finite difference scheme to extract particle velocities $\boldsymbol{v}_{i}$ from positions $\boldsymbol{x}_{i}$.
For a physics-based model, one would need two initial conditions (usually, the current positions and velocities) to predict the next position of the particles. Since we use a first-order finite difference scheme to compute the velocities from positions, this translates to a sequence of two positions of all particles (i.e., $\{\boldsymbol{x}(t-2\Delta t),\boldsymbol{x}(t-\Delta t)\}\to\boldsymbol{x}(t)$). Indeed, this information is sufficient for the GNN to predict the particle trajectories as well. However, longer sequences provide significantly better predictions [10]. Unless stated otherwise, for the results presented here, we have used a sequence of $C=6$ last positions to predict the next position. Note that, therefore, a single full-length computational or experimental positional data with $M$ time-steps results in $(M-6)$ training trajectories (or ‘examples’).
II.2 Node & Edge Features
We closely follow the work of Sanchez-Gonzalez et al. [10] to assign the node and edge features with some key differences that allows an accurate capture of particle size, as well as particle-particle and particle-wall interactions.
Node Features. We use the last $C-1$ velocities of each particle $\boldsymbol{v}_{i}$ ($(C-1)\times\mathrm{DIM}$ values, where $\mathrm{DIM}$ denotes the spatial dimension of the system), its distance to boundaries (shown as $\ell_{i}$’s in Fig. 2; $2\times\mathrm{DIM}$ values), and its radius $a_{i}$ as node features. We also normalize the velocities and distances corresponding to each particle by its radius. We acknowledge that a more physically appropriate normalization would be a constant length scale for all particles. But our scaling, along with what we use for the edge features (see next paragraph), helps the GNN to accurately learn the important characteristics of the system (e.g., clustering).
Edge Features. We consider relative displacement, $\boldsymbol{x}_{ij}=\boldsymbol{x}_{i}-\boldsymbol{x}_{j}$, and absolute distance, $\left\|\boldsymbol{x}_{ij}\right\|$, between particles (see Fig. 2; $\mathrm{DIM}+1$ values), last $C-1$ relative velocities $\boldsymbol{v}_{i}-\boldsymbol{v}_{j}$ ($(C-1)\times\mathrm{DIM}$ values), and sum of the particle radii for each particle pair, $a_{i}+a_{j}$, ($1$ value) as the edge features. We normalize all lengths by $a_{i}+a_{j}$.
Figure 2: Node and edge features of the GNN.
II.3 GNN Architecture
Our GNN architecture is largely similar to that developed by Sanchez-Gonzalez et al. [10], and in particular its PyTorch implementation by Kumar and Vantassel [17]. An outline of our network architecture is provided in Fig. 3. The node and edge feature vectors ($\boldsymbol{v}_{\mathrm{node}}$ and $\boldsymbol{v}_{\mathrm{edge}}$, respectively) get encoded by two Multilayer Perceptrons (MLPs) into their corresponding embeddings $\boldsymbol{\chi}\in\mathbb{R}^{N\times n_{\mathrm{node}}}$ and $\boldsymbol{e}\in\mathbb{R}^{E\times n_{\mathrm{edge}}}$, respectively. Here, symbols stand for number of nodes, $N$; number of edges, $E$; node embedding size, $n_{\mathrm{node}}$; and edge embedding size, $n_{\mathrm{edge}}$. The node and edge embedding sizes are both $128$ and kept the same throughout the network. In each step of the message passing from a neighbor node $j$ to a target node $i$, we concatenate $\boldsymbol{\chi}_{i}$, $\boldsymbol{\chi}_{j}$, and $\boldsymbol{e}_{ij}$ and pass it through an MLP. Then we aggregate all messages from the neighbors of the target node $i$ (i.e, $j\in\mathcal{N}(i)$) to find the overall message $\tilde{\boldsymbol{\chi}}\in\mathbb{R}^{N\times n_{\mathrm{edge}}}$. Note that aggregating all messages is inspired by the physics of the problem; all neighboring particles exert force on the target particle, and the overall force is the summation of all of those interaction forces. Following prior work, we finally apply a residual connection to update the node and edge embeddings. The message passing is performed for $K$ steps. The obtained node embedding $\{\boldsymbol{\chi}\}^{K}$ is then passed through a final MLP to predict the particle accelerations $\hat{\ddot{\boldsymbol{x}}}\in\mathbb{R}^{N\times\mathrm{DIM}}$.
The objective of the model is to predict the acceleration of the particles, given a sequence of $C$ last positions. The loss function to minimize during training is
$\mathcal{L}=\frac{1}{N}\left\|\ddot{\boldsymbol{x}}-\hat{\ddot{\boldsymbol{x}}}\right\|_{2}^{2},$ (1)
where $\ddot{\boldsymbol{x}}$ and $\hat{\ddot{\boldsymbol{x}}}$ are the true and predicted accelerations, and $N$ is the number of particles in example. Following Sanchez-Gonzalez et al. [10], we corrupt the sequence of particle positions with a random walk noise, and use finite difference to compute the accelerations from true positions. For rollout predictions, we use the same finite difference stencil to compute the particle positions form the predicted accelerations.
We have not performed an extensive optimization to fine-tune the network parameters such as the number of message passing steps, number of hidden layers and neurons, etc. However, we found the parameters that work sufficiently well for our objectives. Please see the default parameter values in github.com/rfjd/GNS-DispersedSystems.
Figure 3: An outline of our network architecture including the encoding, message passing, and decoding steps.
III Case Study
Here we consider a case study to put our GNN into test: a multidisperse suspension of spherical disks interacting and sedimenting under gravity in a two-dimensional box $x,y\in[-L,L]$. We perform computer simulations to generate the training and test data. We stress that the following case study is not meant to be a physically accurate representation of the system. However, it has the key characteristics to challenge our GNN model, such as particles of different sizes, intricate interactions between the particles, and physical boundaries. Alternatively, one could collect actual experimental data or some simulation data from a black-box multi-physics simulator to train the model.
III.1 Governing equations
The deterministic dynamics of the particles are governed by the Newton’s second law
$\pi a_{i}^{2}s\rho\frac{d\boldsymbol{v}_{i}}{dt}$ $=-ka_{i}\boldsymbol{v}_{i}+\boldsymbol{F}_{i}^{\text{external}}(\boldsymbol{x})+\boldsymbol{F}_{i}^{\text{interaction}}(\boldsymbol{x}),$ (2a)
$\frac{d\boldsymbol{x}_{i}}{dt}$ $=\boldsymbol{v}_{i}.$ (2b)
Here the symbols stand for the velocity of the particles, $\boldsymbol{v}=\begin{bmatrix}\boldsymbol{v}_{1},\boldsymbol{v}_{2},\dots\end{bmatrix}^{{}^{\intercal}}\in\mathbb{R}^{N\times 2}$; position of the particles, $\boldsymbol{x}=\begin{bmatrix}\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots\end{bmatrix}^{{}^{\intercal}}\in\mathbb{R}^{N\times 2}$; vectors of, respectively, external, $\boldsymbol{F}^{\text{external}}(\boldsymbol{x})\in\mathbb{R}^{N\times 2}$, and interaction, $\boldsymbol{F}^{\text{interaction}}(\boldsymbol{x})\in\mathbb{R}^{N\times 2}$, forces; thickness, $s$, and density, $\rho$, of the particles; and time, $t$. For this ‘toy’ problem, we assume that the drag force on each particle is proportional to the particle’s velocity and radius with a constant of proportionality $k$.
For simplicity, we suppose that the external forces are constant and due to gravity only,
$\boldsymbol{F}_{i}^{\text{external}}=\begin{bmatrix}0&-\pi a_{i}^{2}s(\rho-\rho_{f})g\end{bmatrix},$ (3)
where $\rho_{f}$ is the density of the medium, and $g$ is the gravitational acceleration. We further consider a modified, truncated, Lennard-Jones (LJ) potential to model the inter-particle interactions. Specifically, for particles $i$ and $j$,
$U(r_{ij})=\begin{cases}U_{\text{LJ}}(r_{m})+\frac{dU_{\mathrm{LJ}}(r_{m})}{dr_{ij}}(r_{ij}-r_{m})&r_{ij}\leq r_{m},\\ U_{\text{LJ}}(r_{ij})&r_{m}<r_{ij},\end{cases}$ (4)
with
$U_{\text{LJ}}(r_{ij})=4U_{0}\left(\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{2p}-\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{p}\right).$ (5)
The symbols stand for the particle-particle distance, $r_{ij}$; cutoff distance, $r_{m}$; and depth of the potential well, $U_{0}$. The lower cutoff distance $r_{m}<\sigma_{ij}$ determines the maximum repulsive force for a given set of parameters. We let $\sigma_{ij}=a_{i}+a_{j}$. The LJ force on particle $i$ due to particle $j$ is
$\boldsymbol{F}^{\text{interaction}}_{ij}(\boldsymbol{x})$ $=-\boldsymbol{\nabla}U(r_{ij})=$
$\begin{cases}-\frac{1}{r_{ij}}\frac{dU_{\mathrm{LJ}}(r_{m})}{dr_{ij}}\boldsymbol{x}_{ij}&r_{ij}\leq r_{m},\\ -\frac{1}{r_{ij}}\frac{dU_{\mathrm{LJ}}(r_{ij})}{dr_{ij}}\boldsymbol{x}_{ij}&r_{m}<r_{ij},\end{cases}$ (6)
with $\boldsymbol{x}_{ij}=\begin{bmatrix}x_{i}-x_{j}&y_{i}-y_{j}\end{bmatrix}$, $r_{ij}=\sqrt{x_{ij}^{2}+y_{ij}^{2}}$, and
$\frac{dU_{\mathrm{LJ}}(r_{ij})}{dr_{ij}}=-\frac{4pU_{0}}{r_{ij}}\left(2\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{2p}-\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{p}\right)$ (7)
Accordingly, the total LJ force experienced by particle $i$ is $\sum_{j\neq i}\boldsymbol{F}^{\text{interaction}}_{ij}(\boldsymbol{x})$.
III.2 Boundary conditions
We impose solid-wall boundary conditions with an energy loss ratio $\gamma\in(0,1]$; i.e., upon each collision to a wall, the kinetic energy of the colliding particle is reduced to $\gamma$ multiplied to its pre-collision value. For simplicity we suppose that this energy loss affects both velocity components equally, so that the post-collision velocity is scaled by $\sqrt{\gamma}$. Furthermore, the direction of the particle motion after collision is determined by assuming a mirror reflection from the wall. If we denote the velocity vectors before and after collision by $\boldsymbol{v}$ and $\boldsymbol{v}^{\prime}$, respectively, then such a reflection can be expressed as
$\boldsymbol{v}^{\prime}=\sqrt{\gamma}\left[\boldsymbol{v}-2\left(\boldsymbol{v}\cdot\hat{\boldsymbol{n}}\right)\hat{\boldsymbol{n}}\right],$ (8)
where $\hat{\boldsymbol{n}}$ is the normal unit vector pointing into the domain.
III.3 Numerical Integration
We use the Euler method to numerically integrate the equation of motion (2). Of course, we could use a more accurate numerical scheme such as RK4. But note that the main objective here is to generate some trajectories to train our GNN and put it into a test, rather than developing a comprehensive, accurate, physics-based model and simulator.
IV Results, Discussion, and Future Work
We used a total of $16$ full-length simulations each with approximately $500$ time-steps, which, for $C=6$, translates to $16\times(500-6)=7904$ training examples. We trained our model over $2$ million steps with a batch size of $2$ examples, i.e., $\approx 500$ epochs. For each simulation, a few hundred particles with randomly selected radii from a prespecified list were placed in the upper half of a box at random positions and with random initial two-dimensional velocities.
Our main target for the model was a rollout loss, $\mathcal{L}_{\text{rollout}}$, defined as the mean squared error between the true and predicted positions (normalized by the box size $L$) across all time steps, particles, and spatial dimensions in a full-length test simulation.
Fig. 4 shows a comparison between the ground-truth and predictions of our trained GNN for a test simulation (see Supplementary Video 1 for an animation). The trained GNN accurately predicts the particle positions over long trajectories and captures important spatial structures such as clusters and relative placements of particles. Additionally, it learns the intricate nature of particle-particle and particle-wall interactions/collisions. In particular, the particles never cross the box boundaries. It is important to note that we are not applying any mask or penalty to achieve this. The model’s accuracy slightly drops as the particles get crowded at the bottom of the box. We argue, however, that more training data will resolve this issue. The model simply has not seen enough trajectories in such relatively more complex configurations.
We stress that the particle-size could be treated as a categorical property, where certain particle sizes (categories) were considered in training. As such, particle radius would act as a particle type. However, we were interested in treating the particle radius as a continuous quantity, and in developing a model that predicts the dynamics of a system with any particle size. In fact, we tested our trained model on a binary system with particle sizes not present in the training data. Fig. 5 and Supplementary Video 2 show the corresponding results. The GNN does an excellent job in predicting the particle trajectories despite not seeing these particles sizes during training. We should note however, that our GNN does not show equally accurate results for systems involving particles with radii far outside the radius range used in training. Similarly, one could use training data that include different fluids, particle materials, particle-fluid interfacial properties, etc. to develop a comprehensive model that captures the impact of various physical properties of the problem.
We briefly discuss a few future works below.
  1. (i)
    Inclusion of Rotation. An exciting future work is to account for the inevitable rotation of the particles in two and three dimensions. For such a problem, orientation and angular velocity of the particles contribute to the node and edge features, and the GNN predicts both the translational and rotational accelerations of the particles during training. Since this problem is inherently more complex, it would put our machine learning framework into a more rigorous test. The training data can be obtained through either experiments and high-speed-imaging, or a simulation software such as Ansys Fluent. However, the former, albeit more appealing to us, faces some challenges in a three-dimensional system (e.g., particles masking each other).
Figure 4: Comparison of the ground-truth (left panel) and the GNN prediction (right panel) for the dynamics of a multidisperse suspension of interacting particles sedimenting under gravity in a square box filled with liquid (a test simulation not used for training). $\mathcal{L}_{\text{rollout}}=6.6\times 10^{-4}$.
Figure 5: Comparison of the ground truth (left panel) and the GNN prediction for a system with particle sizes not present during training. $\mathcal{L}_{\text{rollout}}=6.5\times 10^{-4}$.
  1. (ii)
    Asymmetric Particles. Our GNN can be extended to systems with non-spherical particles. One can represent particles with multiple geometrical characteristics instead of a mere radius that we used here for spheres (e.g., length and thickness for a rod-like particle). Additionally, relative orientation of the particles plays an important role in particle-particle interactions/collisions, and particle-wall collisions as well. While we acknowledge that this is a significantly more complex problem, we believe it is a feasible future project.
  2. (iii)
    Colloidal Dispersions. Another avenue for future work is the application of the proposed GNN framework to experimental colloidal dispersions. Recent advances, such as those demonstrated by ActiveNet [18], highlight the potential of GNNs to uncover active and two-body forces in colloidal systems by learning directly from experimental trajectories. Extending our approach to incorporate these techniques could enable the study of forces and dynamics in a wider range of active and passive colloidal systems, including systems under external fields or exhibiting many-body interactions.
The work presented here introduces a new ML framework, a graph neural network (GNN), for a data-driven dynamic simulation of multidisperse suspensions and emulsions in liquids. Our GNN accurately captures the most subtle characteristics of such complex systems, including clustering, as well as particle-particle and particle-wall interactions. This work opens a new avenue of research in using GNNs to simulate dispersed systems.